If a swimming pool has a perimeter of 60 meters and the length is 4 meters longer than the width what is the width?
Answer by oberobic(2304)

Assuming the pool is rectangular, then we know:
Perimeter = 2(Length + Width)
P = 2L + 2W
The perimeter is 60.
P = 60
.
The length is 4 more than the width.
L = W+4
.
2L + 2W = P
Substitute L = W+4 & P = 60
2(W+4) +2W = 60
2W + 8 +2W = 60
Subtract 8 from both sides & collect terms.
4W = 52
Divide both sides by 4
W = 13
.
Substitute this value to find L
L = W+4
L = 13+4
L = 17
.
Always check your work.
2L + 2W = 60 ??
2(17) + 2(13) = 34 + 26
34 + 26 = 60
Check.
.
Reread the question to make sure you answer it.
Answer: The width is 13 meters.
.
Done.